x+y+z=5
-2x-3y+2z=8
3x-y-2z=3

Answer by mananth(16946)

1 x + 1 y + 1 z = 5
3 x -1 y -2 z = 3
-3 x -1 y -2 z = 3

consider equation 1 &2 Eliminate y
Multiply 1 by 1 -5
Multiply 2 by 1 4
we get
1 x + 1 y + 1 z =
3 x -1 y -2 z =
Add the two
4 x 0 y -1 z =
consider equation 2 & 3 Eliminate y
Multiply 2 by -1
Multiply 3 by 1
we get
-3 x + 1 y + 2 z = -3
-3 x -1 y -2 z = 3
Add the two
-6 x + 0 y + 0 z = 0
Consider (4) & (5) Eliminate x
Multiply 4 by 6
Multiply (5) by 4
we get
24 x -6 z = 48
-24 x 0 z = 0
Add the two
0 x -6 z = 48
/ -6
z = -8
-6 x + 0 z= 0
-6 x + -0 = 0
x = 0
Plug x & z in eq (1)
1 x + 1 y + 1 z = 5
-0 + 1 y -8 = 5
1 y = 13
y = 13
x= 0 y= 13 z= -8